The probability of having a sibling is 80%. For 6 randomly chosen people, what is the probability that 5 OR MORE of them have a sibling? Give your answer as 0.XX
Answer #1

probability = 80/100 = 0.80

number of trials n = 6

we have to find P(X is 5 or more) = P(X=5) + P(X=6)

use binomial formula